The Wall Street Journal reported the following spot and forward rates for the Swiss franc ($/SF).

Spot $ 0.8215
30-day forward $ 0.8530
90-day forward $ 0.8553
180-day forward $ 0.8600

a. Was the Swiss franc selling at a discount or premium in the forward market?

  • Discount

  • Premium

b. What was the 30-day forward premium (or discount) percentage? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

c. What was the 90-day forward premium (or discount) percentage? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

d. Suppose you executed a 90-day forward contract to exchange 230,000 Swiss francs into U.S. dollars. How many dollars would you get 90 days hence?

e. Assume a Swiss bank entered into a 180-day forward contract with Bankers Trust to buy $230,000. How many francs will the Swiss bank deliver in six months to get the U.S. dollars? (Round your answer to 2 decimal places.)
Answer #1

a. It is selling at a premium

b. 30-day forward premium (annualized) = ( 0.8530 - 0.8215) / 0.8215 *12 /1 = 46.01%

c. 90-day forward premium (annualized) = ( 0.8553- 0.8215) / 0.8215 *12 /3 = 16.46%

d. dollars would you get 90 days hence = 230,000 * 0.8553 = 196,719

e. francs will the Swiss bank deliver in six months to get the U.S. dollars = 230,000/ 0.8600 = 267,441.86